# On Phase Unwrapping via Digital Wavefront Sensors

### v2

@@ Abstract
In this paper, we consider a new class of methods for phase unwrapping of 2D images based on digital wavefront sensors. Mathematically, many wavefront sensors produce the same measurements of incoming wavefronts regardless of whether they are wrapped or not. Since typical reconstructors for these sensors are optimized to compute smooth wavefronts, it is possible to digitally “propagate” a wrapped phase through such a sensor, resulting in a smooth unwrapped phase. First, we show how this principle can be applied for phase unwrapping using digital Shack-Hartmann and Fourier-type wavefront sensors. Then, we apply our methods to an unwrapping problem appearing in a real-world adaptive optics project currently under development, and compare the results to those obtained with other state-of-the-art algorithms.
@@ 1 Introduction
where the wrapped phase $\phi_{w}$ and the unknown phase $\phi$ are real-valued scalar functions defined for all $\boldsymbol{\xi}\in\Omega\subseteq{\mathbb{R}^{2}}$; see Figure 1.1 for an example on a disc-shaped domain $\Omega$. For this to be a meaningful problem, one typically searches for smooth reconstructions of $\phi$, in particular without any (or only few) $2\pi$ jumps. The phase unwrapping problem (1.1) has applications in numerous fields such as astronomy, biology, or medical imaging, where measurements of the phase of an electric field are often only available in wrapped form. This is because phase information commonly appears in the exponential form $e^{i\phi}$, from which $\phi$ is only determined up to $2\pi$ jumps. Hence, the phase unwrapping problem (1.1) can also be equivalently rephrased as
@@
Over the years, many different methods have been developed to solve the phase unwrapping problem; see e.g. [13, 48, 23, 12, 24]. These work either in the continuous setting of (1.1) and (1.2), or on the actual pixel-discretization of a given wrapped phase, in both cases with the goal to obtain smooth reconstructions without $2\pi$ jumps. In general, one distinguishes between local and global phase unwrapping algorithms. Local methods aim to correct the phase gradient locally, i.e., they follow a certain path and change absolute jumps greater than $\pi$ to their $2\pi$ complement. Such algorithms mainly differ in the choice of the locations where the corrections are applied. A simple strategy is used by the MATLAB built-in function, which unwraps along the first non-singleton dimension of the phase. More advanced algorithms, such as the one described in [18], define a reliability function by using a-priori assumptions about the solution and unwrap the most reliable pixels. Besides local methods, there also exist global phase unwrapping strategies, which aim to reconstruct a phase which satisfies certain a-priori criteria such as smoothness. These are for example based on solving the discrete Poisson equation (PE) [14, 15, 55], or, more recently, on the transport of intensity equation (TIE) [32, 56, 30].
@@
In this paper, we consider a new class of phase unwrapping methods, which are based on mathematical observations of the working principles of certain types of wavefront sensors (WFSs). These sensors are typically used to measure optical wavefront aberrations, and are an integral part of adaptive optics (AO) systems used e.g. in astronomy or medical imaging. Within the commonly used framework of Fourier optics, wavefront aberrations are usually modelled as exponential factors of the form $e^{i\phi}$ in the Fourier plane of the optical system. Using various optical elements, such as lenslet arrays or pyramidal prisms, wavefront sensors obtain implicit measurements of $\phi$ (e.g. its gradient). Using these measurements the full wavefront aberration $\phi$ is then computed via suitable reconstruction algorithms. Crucially, since wavefront aberrations only enter in these sensors via the exponential factor $e^{i\phi}$, the obtained measurements are the same independently of whether $\phi$ contains $2\pi$ jumps or not. This leads directly to the following key idea for phase unwrapping: One can take a wrapped phase $\phi_{w}$ and digitally propagate it through a suitable wavefront sensor, using standard tools of geometrical optics. This provides virtual wavefront measurements, from which one computes $\phi$ using any reconstruction algorithm for the corresponding wavefront sensor. Since these algorithms are designed to produce smooth wavefront reconstructions, the result of this procedure will then be an unwrapped phase $\phi$. The fact that these algorithms are typically optimized with respect to speed and reconstruction quality for aberrations with certain properties then also directly translates to wrapped phases with similar properties.
@@
The aim of this paper is to present the idea of phase unwrapping using digital wavefront sensors outlined above in a detailed and mathematically rigorous way. In particular, we focus on phase unwrapping using digital Shack-Hartmann and Fourier-type wavefront sensors, which are commonly employed in adaptive optics both in astronomy and ophthalmology. The latter is a very broad class of sensors, and includes for example the well-known 3- and 4-sided pyramid wavefront sensors (PWFS). Both the Shack-Hartmann and Fourier-type wavefront sensors have been successfully used for reconstructing wavefront aberrations resulting from atmospheric turbulence, and thus their digital versions should be well suited for unwrapping atmospheric turbulence profiles. This is relevant e.g. in free-space optical communications (FSOC), where such turbulence profiles need to be unwrapped repeatedly and accurately. Hence, the numerical experiments presented in this paper will have a strong focus on this relevant application problem. In particular, we will see that our approach matches, and in some cases outperforms, the reconstruction quality of currently used state-of-the-art methods, and this at a comparable computational cost.
@@ 2 Physical background on wavefront sensors
In this section, we review some necessary background on WFSs, which are a key component of current and upcoming AO systems. In general, AO systems consist of one or more WFSs and deformable mirrors (DMs), and are designed to perform real-time corrections of dynamic wavefront aberrations. The aberrations are corrected by converting wavefront measurements obtained by the WFSs into DM actuator commands such that after reflection on the DMs, an incoming wavefront is restored to its non-aberrated (typically planar) form. For the determination of optimal DM actuator commands, precise phase measurements across the entire wavefront are essential. Hence, reconstructing a wavefront from WFS data and calculating DM actuator commands are two crucial inverse problem.
@@ 2.1 The Shack-Hartmann wavefront sensor
Figure 2.1: Schematic depiction of the physical principle of a SH-WFS. Left and center: Cuts parallel to the optical axis for unperturbed and perturbed wavefronts. Right: Front view onto the sensor array and blurred/shifted images of the distant point source. See Section 2.1 for details. Image taken from [19].
The Shack-Hartmann wavefront sensor (SH-WFS) is one of the most commonly used WFSs in AO [6, 33, 34]. It consists of a quadratic array of small lenses, also called lenslets. Through these incoming light has to pass before reaching a CCD photon detector located in their focal plane. The detector itself is subdivided into a regular grid of so-called subapertures, and each lenslet focuses a part of the incoming light onto one of the subapertures; cf. Figure 2.1. Now assume that the incoming light originates from a distant point source such as an astronomical or artificial star, or a distant laser, and that no aberrations are present. In this case, the incoming wavefront of the light source is plane, and thus each lenslet focuses the light onto a (diffraction-limited) point onto its corresponding subaperture on the CCD detector; cf. Figure 2.1(left). On the other hand, if due to aberrations the wavefront is not plane, then these points are in fact shifted away from the center, and their shape is blurred; cf. Figure 2.1(middle). The SH-WFS effectively measures the change in position of the centers of mass of these blurred points with respect to the centers of the subapertures; cf. Figure 2.1(right). These relative shifts can mathematically be connected to the average gradient (slope) of the wavefront aberration on each subaperture, from which an approximation of the incoming wavefront can then be computed numerically [39]. This is a non-trivial inverse problem, for which many different reconstruction approaches have been proposed; see e.g. [57, 43, 42]. These typically leverage physical priors such as smoothness or statistical assumptions on the wavefront aberrations, and are usually optimized in terms of numerical efficiency to satisfy stringent real-time requirements [46, 49].
@@ 2.2 Fourier-type wavefront sensors
The key component of a Fourier-type WFS is the optical element in its focal plane. Different realisations of Fourier-type WFSs correspond to different optical elements which split the light in the focal plane, e.g. an $n$-sided pyramidal prism. The most common example in AO instruments is the $4$-sided glass pyramidal prism, typically referred to as the pyramid WFS (PWFS) [35, 36]. The prism splits the electromagnetic field into $4$ beams that are imaged in the conjugate pupil plane. Moreover, $2$-sided (roof WFS), $3$-sided, and $6$-sided prisms, as well as the extension of the idea to an infinite number of faces (cone WFS) have been investigated in more detail for astronomical AO [4, 5, 7, 52]. A Fourier-type WFSs for which the optical element does not have the shape of a prism is the Zernike WFS. It has a small circular hole in the center of the mask of the optical element [53]. The iQuad WFS is another Fourier-type WFS. Its focal plane is divided into $4$-quadrants around the origin, i.e., it has a Cartesian structure where each quadrant is phase shifted with its two neighbours [9]. As we will discuss in Section 3.4, the different optical elements can mathematically be characterized via their optical transfer functions (OTFs) [10], examples of which are given in Table 3.1 below. Visualizations of the OTFs corresponding to the 3-sided WFS, cone WFS, iQuad WFS, and different PWFSs and roof WFSs are depicted in Figure 2.3. Note that the apex angle of the prism influences the separation distance between the intensity patterns in the conjugate pupil plane (see Figure 2.3, first three columns from the left).
@@ 3.1 Fourier transforms and Fourier optics
for all (square) integrable functions $f:{\mathbb{R}^{2}}\to\mathbb{C}$. Next, assume that the complex valued electric field $E:{\mathbb{R}^{2}}\to\mathbb{C}$, initially located at the focal length $f$ from the first lens, enters the optical system from the left as depicted in Figure 3.1. Furthermore, assume that in the Fourier plane the resulting field is modulated by the multiplicative factor $e^{i\phi}$, where the function $\phi:{\mathbb{R}^{2}}\to\mathbb{R}$ represents either some wavefront aberration or (the shape of) an optical element such as a pyramidal prism or a spatial light modulator. According to Fourier optics, the electric field $E^{m}:{\mathbb{R}^{2}}\to\mathbb{C}$ exiting the optical system at the focal length $f$ after the second lens can be computed by
@@
$I^{m}(\boldsymbol{x},\phi)=\left|E^{m}(\boldsymbol{x},\phi)\right|^{2}=\left|\mathcal{F}^{-1}\left(\chi_{\Omega}e^{i\phi}\right)(\boldsymbol{x})\right|^{2}\,,\qquad\forall\,\boldsymbol{x}\in\mathbb{R}^{2}\,.$ (3.3)
@@ 3.2 Key ideas and generic unwrapping algorithm
$e^{i\phi(\boldsymbol{\xi})}=e^{i\phi_{w}(\boldsymbol{\xi})}\,,\qquad\forall\,\boldsymbol{\xi}\in\Omega\subset\mathbb{R}^{2}\,,$
and thus by their definition (3.2), there follows the equivalence of the electric fields
@@
This means that when the phase $\phi$ and its wrapped form $\phi_{w}$ are considered as wavefront aberrations in the optical system corresponding to Figure 3.1, their effect on the outgoing electric field $E^{m}$ is equivalent. Consequently, any WFS using the electric field $E^{m}$ as an input for measuring the wavefront aberration is unable to distinguish between the non-wrapped and wrapped phases $\phi$ and $\phi_{w}$, respectively. Crucially, though, most wavefront reconstructors are optimized for reconstructing aberrations with certain smoothness properties, in particular those designed for the SH and Fourier-type WFSs in astronomical applications. Hence, the result of applying such a reconstructor to $E^{m}(\boldsymbol{x},\phi_{w})$ essentially yields a smooth unwrapped phase $\phi_{r}$. Additionally, the closer the properties of $\phi$ conform to those underlying a given WFS reconstruction algorithm, the better an approximation of $\phi$ the unwrapped phase $\phi_{r}$ will be.
From here, only two further observations are necessary. The first one is that WFSs are typically based on the assumption that the incoming light stems from a sufficiently distant point source, corresponding to $E\equiv\delta_{0}$ or $\mathcal{F}(E)\equiv 1$. This is necessary, since otherwise in (3.2) the (non-zero) phase of $\mathcal{F}(E)$ mixes with the aberration $\phi$, making a subsequent separation of the two difficult [19]. Secondly, an electric field $E^{m}$ aberrated by a given wrapped phase $\phi_{w}$ can be obtained completely digitally. For example, in the case of a distant point source one can simply define, cf. (3.3),
@@
Combining these key ideas and observations, we can thus formulate the following generic phase unwrapping algorithm, which will be our basis for further considerations.

### v4

@@ Abstract
In this paper, we derive a new class of methods for the classic 2D phase unwrapping problem of recovering a phase function from its wrapped form. For this, we consider the wrapped phase as a wavefront aberration in an optical system, and use reconstruction methods for (digital) wavefront sensors for its recovery. The key idea is that mathematically, common wavefront sensors are insensitive to whether an incoming wavefront is wrapped or not. However, typical reconstructors for these sensors are optimized to compute smooth wavefronts. Thus, digitally “propagating” a wrapped phase through such a sensor and then applying one of these reconstructors results in a smooth unwrapped phase. First, we show how this principle can be applied to derive phase unwrapping algorithms based on digital Shack-Hartmann and Fourier-type wavefront sensors. Then, we numerically test our approach on an unwrapping problem appearing in a free-space optical communications project currently under development, and compare the results to those obtained with other state-of-the-art algorithms.
@@ 1 Introduction
where the wrapped phase $\phi_{w}$ and the unknown phase $\phi$ are real-valued scalar functions defined for all $\boldsymbol{\xi}\in\Omega\subseteq{\mathbb{R}^{2}}$; see Figure 1.1 for an example on a disc-shaped domain $\Omega$. For this to be a meaningful problem, one typically aims to compute (reconstruct) a smooth unwrapped phase $\phi_{r}$ from the given wrapped phase $\phi_{w}$ as an approximation to the original non-wrapped phase $\phi$, in particular without any (or only few) $2\pi$ jumps. The phase unwrapping problem (1.1) has applications in numerous fields such as astronomy, biology, or medical imaging, where measurements of the phase of an electric field are often only available in wrapped form. This is because phase information commonly appears in the exponential form $e^{i\phi}$, from which $\phi$ is only determined up to $2\pi$ jumps. Hence, the phase unwrapping problem (1.1) can also be equivalently rephrased as
@@
Over the years, many different methods have been developed to solve the phase unwrapping problem; see e.g. [15, 50, 25, 14, 26, 21, 55, 60]. These methods work either in the continuous setting of (1.1) and (1.2), or on the actual pixel-discretization of a given wrapped phase, in both cases with the goal to obtain smooth reconstructions without $2\pi$ jumps. In general, one distinguishes between local and global phase unwrapping algorithms. Local methods aim to correct the phase gradient locally, i.e., they follow a certain path and change absolute jumps greater than $\pi$ to their $2\pi$ complement. Such algorithms mainly differ in the choice of the locations where the corrections are applied. A simple strategy is used by the MATLAB built-in function unwrap(), which unwraps along the first non-singleton dimension of the phase. More advanced algorithms, such as the one described in [20], define a reliability function by using a-priori assumptions about the solution, and unwrap the most reliable pixels. Besides local methods, there also exist global phase unwrapping strategies, which aim to reconstruct a phase which satisfies certain a-priori criteria such as smoothness. These are for example based on solving the discrete Poisson equation (PE) [16, 17, 63], or, more recently, on the transport of intensity equation (TIE) [34, 64, 32]. In recent years, phase unwrapping methods based on deep learning have become popular, see e.g. [21, 55, 60]. For a more detailed discussion of different phase unwrapping methodologies we refer to Section 4.1.
@@
In this paper, we consider a new class of phase unwrapping methods, which are based on mathematical observations of the working principles of certain types of wavefront sensors (WFSs). These sensors are typically used to measure optical wavefront aberrations, and are an integral part of adaptive optics (AO) systems used e.g. in astronomy or medical imaging [42, 35]. Within the commonly used framework of Fourier optics, wavefront aberrations $\phi$ usually enter models of imaging systems in exponential factors of the form $e^{i\phi}$ in the Fourier plane of the optical system. Using various optical elements, such as lenslet arrays or pyramidal prisms, wavefront sensors obtain implicit measurements of $\phi$ (e.g. its gradient $\nabla_{\boldsymbol{\xi}}\,\phi$). Using these measurements, the wavefront aberration $\phi$ is then computed via suitable reconstruction algorithms. Crucially, since wavefront aberrations only enter in these sensors via the exponential factor $e^{i\phi}$, the obtained measurements are the same independently of whether $\phi$ contains $2\pi$ jumps or not. This leads directly to the following key idea for phase unwrapping: One can take a wrapped phase $\phi_{w}$ and digitally propagate it through a suitable wavefront sensor, using standard tools of Fourier optics. This provides virtual wavefront measurements, from which one computes $\phi$ using any reconstruction algorithm for the corresponding wavefront sensor. Since these algorithms are designed to produce smooth wavefront reconstructions, the result of this procedure will then be an unwrapped phase $\phi$. The fact that these algorithms are typically optimized with respect to speed and reconstruction quality for aberrations with certain properties then also directly translates to wrapped phases with similar properties.
@@
The aim of this paper is to present the idea of phase unwrapping using digital wavefront sensors outlined above in a detailed and mathematically rigorous way. In particular, we focus on phase unwrapping using digital Shack-Hartmann and Fourier-type wavefront sensors, which are commonly employed in adaptive optics both in astronomy and ophthalmology. The latter is a very broad class of sensors, and includes for example the well-known 3- and 4-sided pyramid wavefront sensors (PWFS). Both the Shack-Hartmann and Fourier-type wavefront sensors have been successfully used for reconstructing wavefront aberrations resulting from atmospheric turbulence, and thus their digital versions should be well suited for unwrapping atmospheric turbulence profiles. This is relevant e.g. in free-space optical communications (FSOC) [4, 5], where such turbulence profiles need to be unwrapped repeatedly and accurately. Hence, the numerical experiments presented in this paper have a strong focus on this relevant application problem. In particular, we shall see that our approach matches, and in some cases outperforms, the reconstruction quality of currently used state-of-the-art methods, and this also at a comparable computational cost.
@@ 2 Physical background on wavefront sensors
In this section, we review some necessary background on WFSs, which are a key component of current and upcoming AO systems. In general, AO systems consist of one or more WFSs and deformable mirrors (DMs), and are designed to perform real-time corrections of dynamic wavefront aberrations. The aberrations are corrected by converting wavefront measurements obtained by the WFSs into DM actuator commands such that after reflection on the DMs, an incoming wavefront is restored to its non-aberrated (typically planar) form. For the determination of optimal DM actuator commands, precise phase measurements across the entire wavefront are essential. Hence, reconstructing a wavefront from WFS data and calculating DM actuator commands are two crucial inverse problem in AO.
@@ 2.1 The Shack-Hartmann wavefront sensor
Figure 2.1: Schematic depiction of the physical principle of a SH-WFS. Left and center: Cross cuts parallel to the optical axis for unperturbed and perturbed wavefronts. Right: Front view onto the sensor array and blurred/shifted images of the distant point source. See Section 2.1 for details. Image taken from [22].
The Shack-Hartmann wavefront sensor (SH-WFS) is one of the most commonly used WFSs in AO [8, 36, 37]. It consists of a quadratic array of small lenses, also called lenslets. Through these, incoming light has to pass before reaching a CCD photon detector located in their focal plane. The detector itself is subdivided into a regular grid of so-called subapertures, and each lenslet focuses a part of the incoming light onto one of the subapertures; cf. Figure 2.1. Now assume that the incoming light originates from a distant point source such as an astronomical or artificial star, or a distant laser, and that no aberrations are present. In this case, the incoming wavefront of the light source is plane, and thus each lenslet focuses the light onto a (diffraction-limited) point onto its corresponding subaperture on the CCD detector; cf. Figure 2.1(left). On the other hand, if due to aberrations the wavefront is not plane, then these points are in fact shifted away from the center, and their shape is blurred; cf. Figure 2.1(middle). The SH-WFS effectively measures the change in position of the centers of mass of these blurred points with respect to the centers of the subapertures; cf. Figure 2.1(right). These relative shifts can mathematically be connected to the average gradient (slope) of the wavefront aberration on each subaperture, from which an approximation of the incoming wavefront can then be computed numerically [42]. This is a non-trivial inverse problem, for which many different reconstruction approaches have been proposed; see e.g. [65, 45, 44]. These typically leverage physical priors such as smoothness or statistical assumptions on the wavefront aberrations, and are usually optimized in terms of numerical efficiency to satisfy stringent real-time requirements [48, 51].
@@ 2.2 Fourier-type wavefront sensors
The key component of a Fourier-type WFS is the optical element in its focal/Fourier plane. Different realisations of Fourier-type WFSs correspond to different optical elements which split the light in the focal plane, e.g. an $n$-sided pyramidal prism. The most common example in AO instruments is the $4$-sided glass pyramidal prism, typically referred to as the pyramid WFS (PWFS) [38, 39]. The prism splits the electromagnetic field into $4$ beams that are imaged in the conjugate pupil plane. Moreover, $2$-sided (roof WFS), $3$-sided, and $6$-sided prisms, as well as the extension of the idea to an infinite number of faces (cone WFS) have been investigated in more detail for astronomical AO [6, 7, 9, 54]. A Fourier-type WFSs for which the optical element does not have the shape of a prism is the Zernike WFS. It has a small circular hole in the center of the mask of the optical element [58]. The iQuad WFS is another Fourier-type WFS. Its focal plane is divided into $4$-quadrants around the origin, i.e., it has a Cartesian structure where each quadrant is phase shifted with its two neighbours [11]. As we will discuss in Section 3.4, the different optical elements can mathematically be characterized via their optical transfer functions (OTFs) [12], examples of which are given in Table 3.1 below. Visualizations of the OTFs corresponding to the 3-sided WFS, cone WFS, iQuad WFS, and different PWFSs and roof WFSs are depicted in Figure 2.3. Note that the apex angle of the prism influences the separation distance between the intensity patterns in the conjugate pupil plane (see Figure 2.3, first three columns from the left).
@@ 3.1 Fourier transforms and Fourier optics
for all (square) integrable functions $f:{\mathbb{R}^{2}}\to\mathbb{C}$. Next, assume that the complex valued electric field $E:{\mathbb{R}^{2}}\to\mathbb{C}$, initially located at the focal length $f$ from the first lens, enters the optical system from the left as depicted in Figure 3.1. Furthermore, assume that in the Fourier plane the resulting field is modulated by the multiplicative factor $e^{i\phi}$, where the function $\phi:{\mathbb{R}^{2}}\to\mathbb{R}$ represents either some wavefront aberration or any other type of phase modulation (e.g., induced by an optical element or a spatial light modulator). According to Fourier optics, the electric field $E^{m}:{\mathbb{R}^{2}}\to\mathbb{C}$ exiting the optical system at the focal length $f$ after the second lens can be computed by
@@
$I^{m}(\boldsymbol{x},\phi)=\left|E^{m}(\boldsymbol{x},\phi)\right|^{2}=\left|\mathcal{F}^{-1}\left(\chi_{\Omega}(\boldsymbol{\xi})e^{i\phi(\boldsymbol{\xi})}\right)(\boldsymbol{x})\right|^{2}\,,\qquad\forall\,\boldsymbol{x}\in\mathbb{R}^{2}\,.$ (3.3)
@@ 3.2 Key ideas and generic unwrapping algorithm
[…]
and thus, by its definition (3.2), there follows the equivalence of the electric fields
@@
This means that when the phase $\phi$ and its wrapped form $\phi_{w}$ are considered as wavefront aberrations in the optical system corresponding to Figure 3.1, their effect on the outgoing electric field $E^{m}$ is equivalent. Consequently, any WFS using the electric field $E^{m}$ as an input for measuring the wavefront aberration is unable to distinguish between the non-wrapped and wrapped phases $\phi$ and $\phi_{w}$, respectively. Crucially, though, most wavefront reconstructors are optimized for reconstructing aberrations with certain smoothness properties, in particular those designed for the SH and Fourier-type WFSs in astronomical applications. Hence, the result of applying such a reconstructor to $E^{m}(\boldsymbol{x},\phi_{w})$ essentially yields a smooth and unwrapped phase, denoted by $\phi_{r}$.
From here, only two further observations are necessary. The first one is that WFSs are typically based on the assumption that the incoming light stems from a sufficiently distant point source, which corresponds to $E\equiv\delta_{0}$ or $\mathcal{F}(E)\equiv 1$. This is necessary, since otherwise in (3.2) the (non-zero) phase of $\mathcal{F}(E)$ mixes with the aberration $\phi$, making a subsequent separation of the two difficult [22]. Secondly, an electric field $E^{m}$ aberrated by a given wrapped phase $\phi_{w}$ can be obtained completely digitally. For example, in the case of a distant point source one can simply define, cf. (3.3),
@@
Combining these key ideas and observations, we can thus formulate the following generic phase unwrapping algorithm, which is the basis for our further considerations.
@@
Remark.
The phase unwrapping problem can be considered both in a continuous and a discrete setting. Our generic Algorithm 1 and its specific instances discussed below are operating in the continuous setting corresponding of problem (1.1). Besides the greater generality of the continuous setting, a main reason for this is that the underlying physical models of the considered (digital) wavefront sensors operate in the continuous setting as well. When our proposed algorithms are applied in the discrete setting of numerical computations, the involved operations need to be replaced by their discrete counterparts. This essentially concerns the continuous Fourier transform, which has to be replaced by the (fast) discrete Fourier transform, as is done in the numerical experiments presented in Section 4. Note that this situation is the same for the global comparison algorithms based on the PE and the TIE, which were introduced above and are discussed in detail in Section 4.1. These methods are formulated in terms of (continuous) partial differential equations, which are discretized in numerical applications.